An electric charge of 2.10-3C at a point X in an electric field had electric potential energy of 4.10-2 J. It was moved to another point Y where it’s potential energy was 6.10-2J. What is the potential difference between X and Y and the electric field strength if the points X and Y and seperated by distance 0.25m?


1
Expert's answer
2021-04-19T17:11:03-0400

Solution:


ΔEPE = 0.0610J - 0.0410J = 0.0200 J, so

a) ΔVxy = "\dfrac{0.02J}{2.1*10^{-3}C}=9.52V"


b) E = "\dfrac{ΔV_{xy}}{d}=\dfrac{9.52V}{0.25m}=38.1(\frac{V}{m})"

assuming the motion was parallel to the field.